(1 point) This problem has two parts. The second part will appear after you answer the first part correctly. (a) What are the vertical asymptotes of f()2 Vertical asymptotes at z5,-5 (b) Find the following left- and right-hand limits at the vertical asymptote 5. Enter I for oo, I for-oo, or DNE for does not exist. Your answer should be a number, a list of numbers separated by commas, or None 25 3z 3r2 lim 25 22-25 Find the following left-and right-hand limits at the vertical asymptote z = 5. 3z2 2 25 lim
Answer #1

b) 3r2 r→-5-12-25 = \infty

\lim_{x\to-5^+}\frac{3x^2}{x^2-25} = -\infty

\lim_{x\to+5^-}\frac{3x^2}{x^2-25} = -\infty

\lim_{x\to+5^+}\frac{3x^2}{x^2-25} = \infty